question 11

question 11

question 11
Answer #1

first we have to find the inpedance in the circuit then

X_{L}=\omega *L=2*\pi*f*l=2*\pi*70*0.0940=41.34\: ohm in +90 degrees

X_{C}=\frac{1}{\omega*C }=\frac{1}{2*\pi*f*C}=\frac{1}{2*\pi*70*25.5}=90.9456 \: ohm in -90 degrees

now we have to find the total impedance Z then

\vec{Z}=R+(X_{L}+X_{C})

\vec{Z}=25+(41.34-90.94)*j=25-(49.60)*j

now we obtain the phese angle then

\theta=tan^{-1}(\frac{-49.60}{25})=-63.25^{0}

now the system is capative because

Xc > XL